# RL4Med-DDPO: Reinforcement Learning for Controlled Guidance Towards Diverse Medical Image Generation using Vision-Language Foundation Models

### v1

@@ Abstract
Vision-Language Foundation Models (VLFM) have shown a tremendous increase in performance in terms of generating high-resolution, photorealistic natural images. While VLFMs show a rich understanding of semantic content across modalities, they often struggle with fine-grained alignment tasks that require precise correspondence between image regions and textual descriptions—a limitation in medical imaging, where accurate localization and detection of clinical features are essential for diagnosis and analysis. To address this issue, we propose a multi-stage architecture where a pre-trained VLFM provides a cursory semantic understanding, while a reinforcement learning (RL) algorithm refines the alignment through an iterative process that optimizes for understanding semantic context. The reward signal is designed to align the semantic information of the text with synthesized images. We demonstrate the effectiveness of our method on a medical imaging skin dataset where the generated images exhibit improved generation quality and alignment with prompt over the fine-tuned Stable Diffusion. We also show that the synthesized samples could be used to improve disease classifier performance for underrepresented subgroups through augmentation.
@@
Medical Image Generation Policy Optimization Reinforcement Learning Vision-Foundation Models.
@@ 1 Introduction
Synthetic image generation is of importance, especially in medical imaging, as it can be used for tasks such as data augmentation [16], debasing classifiers [17] or accurate detection and diagnosis of disease [11, 29]. Additionally, high-resolution and precise image-generation capabilities in complex settings, such as drug discovery or personalized diagnosis, require analyzing counterfactual "what if" scenarios [20, 18]. Recently, Denoising Diffusion models (DDMs) [13] have shown outstanding performances in high-resolution conditional image generation. Despite their impressive capability to synthesize images, they are prone to biases. These diffusion architectures utilize a controlled sampling process, which can either be classifier-free [14] or classifier-guided [7]. A promising alternative recently proposed for this goal is the use of Reinforcement Learning (RL) to optimize the diffusion process for improved control and adaptability [2, 9, 19, 31]. Fine-tuning the diffusion model to optimize a desired reward function can enable these models to incorporate task-specific preferences, potentially reducing bias and improving alignment between generated samples and predefined constraints. Denoising Diffusion Policy Optimization (DDPO) [2] is an RL-based method that reframes the diffusion process as a multi-step Markov Decision Process (MDP) to optimize a given reward function. With the rise of policy-based methods, the effectiveness of various reward functions has been demonstrated in natural imaging domains, including diversity-based rewards [31, 19], alignment rewards [2], and visual rewards such as aesthetic quality [2]. However their applicability in medical imaging remains underexplored.
@@ 2.2 Training Details
Finetuning Stable Diffusion: The original Stable Diffusion v1.5 [22] architecture consists of four main components: (1) image encoder, (2) Contrastive Language-Image Pre-training (CLIP) text encoder, (3) denoising U-Net for reverse diffusion and (4) an image decoder. Aside from the denoising U-Net, all other components of the Stable diffusion model remain frozen during fine-tuning. The input to the CLIP encoder is a template text: A dermoscopic image with [disease] showing [artifact], where disease is melanoma or melanocytic nevis and artifact is hairs, gel bubbles, ink or ruler thus helping to synthesize semantically aligned images from the input text.
@@ 3.3 Quantitative Evaluations
# Real Sample 43 260 4 440 3051 489 21 520
APR $\uparrow$ (%) SD 9.20 27.00 3.70 0.60 1.20 15.00 1.00 0.00
SD+RL 29.80 55.70 43.40 0.90 41.50 28.60 33.50 0.60
Table 3: Quantitative results for evaluating synthesized samples for 32 different prompts (1000 samples per prompt) of various artifacts with MEL or NV. LPIPS was calculated by randomly selecting 1000 samples per prompt and comparing them sequentially with corresponding real images, resulting in a total of 32,000 comparisons.

### v2

# equal contribution

@@ Abstract
Vision-Language Foundation Models (VLFM) have shown a tremendous increase in performance in terms of generating high-resolution, photorealistic natural images. While VLFMs show a rich understanding of semantic content across modalities, they often struggle with fine-grained alignment tasks that require precise correspondence between image regions and textual descriptions, a limitation in medical imaging, where accurate localization and detection of clinical features are essential for diagnosis and analysis. To address this issue, we propose a multi-stage architecture where a pre-trained VLFM (e.g. Stable Diffusion) provides a cursory semantic understanding, while a reinforcement learning (RL) algorithm refines the alignment through an iterative process that optimizes for understanding semantic context. The reward signal is designed to align the semantic information of the text with synthesized images. Experiments on the public ISIC2019 skin lesion dataset demonstrate that the proposed method improves (a) the quality of the generated images, and (b) the alignment with the text prompt over the original fine-tuned Stable Diffusion baseline. We also show that the synthesized samples could be used to improve disease classifier performance for underrepresented subgroups through augmentation. Our code is accessible through the project website. [^1] https://parhamsaremi.github.io/rl4med-ddpo
@@
Medical Image Generation Policy Optimization Reinforcement Learning Vision-Language Foundation Models.
@@ 1 Introduction
Synthetic image generation is of importance, especially in medical imaging, as it can be used for tasks such as data augmentation [17], debasing classifiers [18] or accurate detection and diagnosis of disease [12, 31]. Additionally, high-resolution and precise image-generation capabilities in complex settings, such as drug discovery or personalized diagnosis, require analyzing counterfactual "what if" scenarios [10, 22, 19]. Recently, Denoising Diffusion models (DDMs) [14] have shown outstanding performances in high-resolution conditional image generation. Despite their impressive capability to synthesize images, they are prone to biases. These diffusion architectures utilize a controlled sampling process, which can either be classifier-free [15] or classifier-guided [7]. A promising alternative recently proposed for this goal is the use of Reinforcement Learning (RL) to optimize the diffusion process for improved control and adaptability [2, 9, 21, 33]. Fine-tuning the diffusion model to optimize a desired reward function can enable these models to incorporate task-specific preferences, potentially reducing bias and improving alignment between generated samples and predefined constraints. Denoising Diffusion Policy Optimization (DDPO) [2] is an RL-based method that reframes the diffusion process as a multi-step Markov Decision Process (MDP) to optimize a given reward function. With the rise of policy-based methods, the effectiveness of various reward functions has been demonstrated in natural imaging domains, including diversity-based rewards [33, 21], alignment rewards [2], and visual rewards such as aesthetic quality [2]. However their applicability in medical imaging remains underexplored.
@@ 2.2 Training Details
Finetuning Stable Diffusion: The original Stable Diffusion v1.5 [24] architecture consists of four main components: (1) image encoder, (2) Contrastive Language-Image Pre-training (CLIP) text encoder, (3) denoising U-Net for reverse diffusion and (4) an image decoder. Aside from the denoising U-Net, all other components of the Stable diffusion model remain frozen during fine-tuning. Similar to PRISM [20], the input to the CLIP encoder is a template text: A dermoscopic image with [disease] showing [artifact], where disease is melanoma or melanocytic nevis and artifact is hairs, gel bubbles, ink or ruler thus helping to synthesize semantically aligned images from the input text.
@@ 3.3 Quantitative Evaluations
# Real Sample 543 260 4 440 3051 489 21 520
APR $\uparrow$ (%) SD (Baseline) 63.74 13.75 0.2 76.56 44.37 13.54 0.1 80
SD+RL 86.97 94.37 1.6 93.85 75.72 80.62 1.4 87.18
Table 3: Quantitative results for evaluating synthesized samples for 32 different prompts (1000 samples per prompt) of various artifacts with MEL or NV. LPIPS was calculated by randomly selecting 1000 samples per prompt and comparing them with corresponding real images, resulting in a total of 32,000 comparisons.